Calculate the frequency of the oscillations of a simple pendulum of length 984 mm. Give your answer to an appropriate number of significant figures


1
Expert's answer
2021-02-23T10:04:58-0500
"T=2\pi\sqrt{\dfrac{L}{g}}=2\pi\sqrt{\dfrac{0.984\ m}{9.8\ \dfrac{m}{s^2}}}=1.989\ s,""f=\dfrac{1}{T}=\dfrac{1}{1.989\ s}=0.503\ Hz."